# Convergence for Natural Policy Gradient on Infinite-State Average-Reward Markov Decision Processes

### v1

@@ Abstract.
Infinite-state Markov Decision Processes (MDPs) are essential in modeling and optimizing a wide variety of engineering problems. In the reinforcement learning (RL) context, a variety of algorithms have been developed to learn and optimize these MDPs. At the heart of many popular policy-gradient based learning algorithms, such as natural actor-critic, TRPO, and PPO, lies the Natural Policy Gradient (NPG) algorithm. Convergence results for these RL algorithms rest on convergence results for the NPG algorithm. However, all existing results on the convergence of the NPG algorithm are limited to finite-state settings.
We prove the first convergence rate bound for the NPG algorithm for infinite-state average-reward MDPs, proving a $O(1/\sqrt{T})$ convergence rate, if the NPG algorithm is initialized with a good initial policy. Moreover, we show that in the context of a large class of queueing MDPs, the MaxWeight policy suffices to satisfy our initial-policy requirement and achieve a $O(1/\sqrt{T})$ convergence rate. Key to our result are state-dependent bounds on the relative value function achieved by the iterate policies of the NPG algorithm.
@@ 1. Introduction
Infinite-state Markov Decision Processes (MDPs) are used to model many engineering problems. A state might include the number of orders for a product, the time since an event occurred, or the number of people waiting in an unbounded queue. While artificial truncation could be imposed, often the cleanest mathematical description of a problem will have infinitely many states.
In various disciplines, people have tried to study and optimize these MDPs using a variety of tools, including methods based off of dynamic programming, such as value iteration and policy iteration methods (Bertsekas, 2012; Puterman, 2014), as well as gradient-based methods drawing on continuous optimization.
In the reinforcement learning (RL) context, algorithms have been developed to simultaneously learn the structure of the MDPs and optimally solve them. While most standard learning theory focuses on the settings with finite state spaces, some recent work has empirically established that learning-based approaches can be used in the infinite-state setting as well (Dai and Gluzman, 2022).
An important and popular class of learning algorithms are policy-gradient-based methods, including the actor-critic algorithm (Sutton and Barto, 1998), the natural actor-critic algorithm (Peters and Schaal, 2008), and more recent methods such as the Trust Region Policy Optimization (TRPO) (Schulman et al., 2015) and Proximal Policy Optimization (PPO) algorithms (Schulman et al., 2017; Dai and Gluzman, 2022).
In this paper, we study the natural policy gradient (NPG) algorithm,which lies at the heart of policy-gradient methods such as natural actor-critic, TRPO, and PPO (Kakade, 2001), Algorithm 1. While the interpretation of the original NPG algorithm is in terms of the gradient of the objective of the MDP with respect to a value function parameterization, it can also be thought as a “softer” version of policy iteration. In this viewpoint, instead of finding the optimal policy with respect to the current estimate of the value function as in policy iteration, we find a randomized approximation to the optimal policy. This encourages exploration of all states and actions, which is important in the RL context. In the RL algorithm, the policy evaluation part of policy iteration is further replaced by a learning algorithm such as TD learning, Using this viewpoint, the convergence of the overall algorithm can be decomposed into an analysis of NPG with perfect policy evaluation and the analysis of TD learning; see, for example, (Abbasi-Yadkori et al., 2019). In this paper, we focus on the analysis of NPG with perfect policy evaluation. Understanding its behavior is critical to understanding the behavior of the overall RL algorithm (Abbasi-Yadkori et al., 2019).
In recent years, progress has been made into understanding the convergence rate of these learning-based algorithms, building upon convergence-rate results for the NPG algorithm. However, existing results on the convergence of the NPG algorithm are limited to settings with finite (but potentially large) state spaces (Murthy and Srikant, 2023; Even-Dar et al., 2009; Agarwal et al., 2021).
We want to investigate the convergence rate of the NPG algorithm in infinite-state, average-reward MDPs, inspired by the MDPs which arise when optimizing queueing systems. Specifically, our goal is to establish theoretical bounds on the convergence rate of the NPG algorithm to the globally optimal policy.
@@
  • In Theorem 1, we prove that for any infinite-state average-reward MDP, given mild assumptions on the structure of the MDP and a good initial policy, the NPG algorithm converges to the global optimum policy. Moreover, we prove an $O(1/\sqrt{T})$ convergence rate bound. This result is the first convergence result for the NPG algorithm in the infinite-state setting. Prior work in the finite-state setting required strong assumptions on all policies and over all states, as we discuss in Section 4.1 (Murthy and Srikant, 2023; Even-Dar et al., 2009). In contrast, our result only makes assumptions on the initial policy.
  • In Theorem 3, we examine a large class MDPs arising from queueing systems, and show that the MaxWeight policy satisfies our quality requirement for the initial policy. Therefore, initializing the NPG algorithm with MaxWeight will lead to an $O(1/\sqrt{T})$ convergence rate.
  • Methodologically, the key step in our proof is to establish a bound the growth rate of the relative value function, relative to the reward available in a given state. We do this by showing that the relative value function only mildly changes between the iterations of NPG, in conjunction with our assumption on the growth rate of the initial policy. Armed with the growth rate bound, we make a novel alteration to the NPG algorithm by setting a corresponding state-dependent step size. We prove that our state-dependent step size adjusts for the growing relative value function, allowing us to prove our convergence result. We discuss our approach in more detail in Section 4.1.
This paper focuses on infinite-state average-reward MDPs, which are the MDPs of most interest in a variety of engineering contexts. While infinite-state discounted-reward MDPs have rarely been explicitly studied, results on the NPG algorithm in such settings follow from existing finite-state NPG results in the discounted setting. This is in contrast to infinite-state average-reward results, for which new techniques were needed. We discuss infinite-state discounted-reward MDPs in Section 7.
@@
  • Section 3: We define the MDP model, give our mild assumptions on the MDP, and define the Natural Policy Gradient algorithm.
  • Section 4: We state our convergence result for the NPG algorithm in the infinite-state, average-reward setting, and discuss the key challenges and a sketch of our proof.
  • Section 5: We prove our convergence result for the NPG in the infinite-state, average-reward setting.
  • Section 6: We demonstrate that our results are applicable to a natural class of infinite-state average-reward MDPs arising out of queueing theory.
@@ 2. Prior work
We overview prior results on the NPG algorithm in the finite-state setting in Section 2.1, connections between the NPG algorithm and a variety of reinforcement learning algorithms in Section 2.2, results on the NPG algorithm in highly-structured infinite-state settings in Section 2.3, and prior work applying reinforcement learning to queueing problems in Section 2.4.
@@ 2.1. Natural Policy Gradient with finite state-space
The Natural Policy Gradient (NPG) algorithm for MDP optimization utilizes the Fisher information associated with a policy to choose a gradient descent direction to update a MDP policy (Kakade, 2001). The NPG algorithm has been shown to have attractive properties, including global convergence in the tabular setting with finite state space (Agarwal et al., 2021).
@@ 2.2. NPG and Reinforcement Learning
The theoretical analysis of the NPG algorithm underpins the theoretical motivation for these reinforcement learning algorithms. We prove the first convergence result for NPG in the infinite-state average-reward setting, giving theoretical backing for the use of NPG-based reinforcement learning algorithms in this setting.
@@ 2.3. Natural Policy Gradient for specialized settings
Fazel et al. (2018) study the Linear Quadratic Regulator (LQR), an important problem in control theory which can be thought of as an MDP optimization problem with an uncountably infinite state-space and average reward. The paper proves that in this setting, the naive policy gradient algorithm converges to the optimal policy, and the NPG algorithm converges with a faster rate guarantee. Kunnumkal and Topaloglu (2008) study the base-stock inventory control problem, and prove that an algorithm using the policy gradient framework achieves convergence to the optimal policy, with a bound on the convergence rate.
Follow-up study of these settings has recently demonstrated that these settings exhibit additional structure which allows these convergence results to hold (Bhandari and Russo, 2019). This additional structure relates to the interplay between policy gradient algorithms, which the above results focus on, and policy improvement algorithms. In particular, it has been shown that if the standard policy improvement algorithm exhibits no suboptimal stationary points, then policy gradient algorithms will track the standard policy iteration algorithm and thereby converge to the global optimum. This mechanism is the root cause behind the prior convergence results in the LQR and base-stock settings.
In contrast, such properties do not hold for general infinite-state MDPs, such as the ones we study, which have much less structure than these previously-studied settings.
@@ 2.4. Queueing and Reinforcement Learning
In this paper, we study infinite-state average-reward MDP optimization, and use MDPs arrising from queueing theory as an example application of our results (See Section 6).
A variety of papers have studied applying reinforcement learning to queueing problems, including using learning algorithms which build off of the NPG algorithm.
Dai and Gluzman (2022) study variance reduction techniques for the PPO algorithm in the context of queueing models, and demonstrate empirically strong performance, converging towards an optimal policy in several queueing settings. Che et al. (2023) further employ a differentiable simulation-based modeling technique, allowing additional approximate gradient information to be derived, which is then used to improve the empirical reinforcement learning performance in a queueing context.
Wei et al. (2023) have demonstrated that a “sample augmentation” technique can reduce the amount of sample data necessary to converge towards an optimal policy for pseudo-stochastic reinforcement learning settings, including queueing models. However, the paper focuses on discounted-cost finite-state queueing MDPs, rather than the infinite-state average-reward queueing models which are preferred elsewhere in queueing theory.
In preprint work, Adler and Subramanian (2023) study an infinite-state, average-reward MDP optimization setting, where the underlying MDP is parameterized by an unknown parameter $\theta$. Queueing models with unknown dynamics are used as the motivating models. The paper assumes that an optimal policy $\pi_{\theta}^{*}$ is known for the MDP under any given parameter $\theta$, and focuses on learning the true parameter value $\theta^{*}$, and on minimizing the regret experienced while learning that parameter. Our result complements this paper, as we show that the NPG algorithm can be used to find optimal policies for MDPs in a setting where the underlying parameters are exactly known, providing the necessary input $\pi_{\theta}^{*}$ for the above result.
In work currently under submission, Chen et al. study a primal-dual optimization technique for optimizing constrained MDPs, such as those encountered in queueing problems. They consider a Langrangian relaxation of the original MDP, and simultaneously optimize the policy and the dual weights of the relaxation. In the special case of an unconstrained MDP, the underlying primal algorithm they study is the NPG algorithm.
@@ 3. Model, Assumptions, and Algorithm
We introduce our MDP model in Section 3.1, our mild assumptions on the structure of the MDP in Section 3.2, and the NPG algorithm in Section 3.3.
@@ 3.1. Markov Decision Processes
The reward is given by a reward function $r(s,a)$ which maps a state-action pair to a real-valued reward. We likewise write $r(s,\pi)$ to denote the expected single-step reward associated with policy $\pi$ and state $s$. We define $r_{\max}(s)$ to be the maximum reward achievable in state $s$, $r_{\max}(s):=\max_{a}r(s,a)$.
@@
$J_{\pi}+V_{\pi}(s)=r(s,\pi)+\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{\pi}(s)}[V_{\pi}(s^{\prime})]$
@@
$J_{\pi}+Q_{\pi}(s,a)$ $=r(s,a)+\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{\pi}(s)}[Q_{\pi}(s^{\prime},\pi)].$
To uniquely specify the additive constant $C$, we will adopt the convention that $V_{\pi}(\vec{0})=0$, for a specially identified state $\vec{0}$ defined in Assumption 1. In the queueing setting discussed in Section 6, $\vec{0}$ is the state with no jobs in the system.
@@
3.2. Assumptions
We make two assumptions on the structure of our MDP: One on the rewards in the MDP, and one on the connectedness of high-reward states. We verify these assumptions for a large class of queueing MDPs in Theorem 3.
First, we make mild assumptions on the reward structure of the MDP:
Assumption 1 (Reward structure).
  1. (a)
  1. (b)
  1. (c)
  2. (d)
Second, we assume that the high-reward states are uniformly connected, under an arbitrary policy:
Assumption 2 (Uniform connectedness of high-reward states).

### v2

# Convergence for Natural Policy Gradient on Infinite-State Queueing MDPs

@@ Abstract.
A wide variety of queueing systems can be naturally modeled as infinite-state Markov Decision Processes (MDPs). In the reinforcement learning (RL) context, a variety of algorithms have been developed to learn and optimize these MDPs. At the heart of many popular policy-gradient based learning algorithms, such as natural actor-critic, TRPO, and PPO, lies the Natural Policy Gradient (NPG) policy optimization algorithm. Convergence results for these RL algorithms rest on convergence results for the NPG algorithm. However, all existing results on the convergence of the NPG algorithm are limited to finite-state settings.
We study a general class of queueing MDPs, and prove a $O(1/\sqrt{T})$ convergence rate for the NPG algorithm, if the NPG algorithm is initialized with the MaxWeight policy. This is the first convergence rate bound for the NPG algorithm for a general class of infinite-state average-reward MDPs. Moreover, our result applies to a beyond the queueing setting to any countably-infinite MDP satisfying certain mild structural assumptions, given a sufficiently good initial policy. Key to our result are state-dependent bounds on the relative value function achieved by the iterate policies of the NPG algorithm.
@@ 1. Introduction
Policy optimization for queueing systems is a major challenge, both for system operators and for theorists. While explicit optimal policies may be known in highly simplified settings or in asymptotic regimes, for general systems algorithmic optimization is required. A popular approach is to use optimization techniques from reinforcement learning (RL) to find optimal control policies for these queueing systems, such as optimal scheduling policies. For instance, the Natural Policy Gradient (NPG) algorithm lies at the heart of many popular RL algorithms, such as natural actor-critic, Trust Region Policy Optimization (TRPO) and Proximal Policies Optimization (PPO). While reinforcement learning techniques are widely used in queueing systems with unbounded queues, theoretical progress has not kept up with practical usage. The convergence of NPG in the setting of countably infinite Markov decision problems MDPs, including the infinite MDPs that arise when optimizing queueing systems, is an open problem.
The NPG algorithm has been well-studied in recent years, but only in settings with finite state spaces (Even-Dar et al., 2009; Murthy and Srikant, 2023; Agarwal et al., 2021). Convergence results for the NPG algorithm in these settings have been proven, and from these results, convergence results for NPG-based reinforcement learning algorithms have also been proven. The standard progression in this area is to start by studying NPG in the perfect-information, tabular setting, and then build on the ensuing results to analyze settings with learning and with function approximation.
We initiate that progression in the setting of countably-infinite queueing MDPs, or more generally countably-infinite MDPs with certain continuity properties and a single, connected high-reward region of the Markov chain, as we discuss in Section 3.5. These properties are satisfied by a wide variety of MDPs that arise in the setting of queueing theory, as well as many MDPs arising from other settings.
In this setting, we prove that the NPG algorithm converges to an optimal policy, with a concrete bound on the convergence rate. Our result is the first convergence result for perfect-information NPG in any general countably-infinite setting.
@@
  • In Theorem 1, we prove that in any queueing system in general class of queueing systems known as the Generalized Switch with Static Environment, the NPG algorithm converges to the globally optimal policy, for the mean queue length objective, or more generally the $\alpha$th-moment of queue length for any $\alpha\geq 1$. Moreover, we prove an $O(1/\sqrt{T})$ convergence rate bound. This result is the first convergence result for the NPG algorithm in any infinite-state queueing setting.
  • More generally, in Theorem 2, we prove that in any countably-infinite average-reward MDP, given a sufficiently good initial policy, the NPG algorithm converges to the optimal policy with an $O(1/\sqrt{T})$ convergence rate bound, given mild structural assumptions on the MDP. This is the first such result for any general class of infinite-state average-reward MDPs. Prior work in the finite-state setting required strong assumptions on all policies and over all states, as we discuss in Section 4.1 (Murthy and Srikant, 2023; Even-Dar et al., 2009).
  • Methodologically, we have two key novelties in our proof. First, we initialize the NPG algorithm with the MaxWeight policy, rather than with an arbitrary policy, as in previous proofs in the finite state-space setting. The low growth rate of the relative value function of MaxWeight in particular is key to our result. Second, we show that the relative value function only mildly changes between iterations of NPG, allowing us to bound the growth rate of the relative value function of all iterations of the NPG algorithm. Using a state-dependent NPG step size, we adjust for the growing relative value function and prove our convergence result. We discuss our approach in more detail in Sections 4.1 and 4.2.
This paper focuses on average-reward performance objectives for queueing problems, rather than discounted-reward objectives, as average-reward objectives are most often of importance to queueing practitioners. While infinite-state discounted-reward MDPs have rarely been explicitly studied, results on the NPG algorithm in such settings follow from existing finite-state NPG results in the discounted setting, given the special structure of queueing MDPs. This is in contrast to infinite-state average-reward results, for which new techniques were needed. We discuss infinite-state discounted-reward MDPs in Section 6.
@@
  • Section 3: We define the MDP model, define our queueing model, and define the Natural Policy Gradient algorithm, and define the key properties of queueing MDPs that allow our proof to work.
  • Section 4: We state our convergence result for the NPG algorithm in the setting of countably-infinite average reward queueing MDPs, and discuss the key challenges and a sketch of our proof.
  • Section 5: We prove our convergence result for the NPG algorithm for queueing MDPs.
@@ 2. Prior work
We overview prior results on the NPG algorithm in the finite-state setting in Section 2.1, connections between the NPG algorithm and a variety of reinforcement learning algorithms in Section 2.2, results on the NPG algorithm in highly-structured infinite-state settings in Section 2.3, and prior work applying reinforcement learning to queueing problems in Section 2.4. Finally, we discuss concurrent work on a related, but distinct setting in Section 2.5.
@@ 2.1. Natural Policy Gradient with finite state-space
The Natural Policy Gradient (NPG) algorithm for MDP optimization utilizes the Fisher information associated with a policy to choose a gradient descent direction to update an MDP policy (Kakade, 2001). The NPG algorithm has been shown to have attractive properties, including global convergence in the tabular setting with finite state space (Agarwal et al., 2021; Even-Dar et al., 2009; Murthy and Srikant, 2023).
@@ 2.2. NPG and Reinforcement Learning
The theoretical analysis of the NPG algorithm underpins the theoretical motivation for these reinforcement learning algorithms. We prove the first convergence result for NPG in the infinite-state average-reward setting, and specifically in the setting of queueing MDPs, giving theoretical backing for the use of NPG-based reinforcement learning algorithms in this setting.
@@ 2.3. Natural Policy Gradient for specialized settings
Fazel et al. (Fazel et al., 2018) study the Linear Quadratic Regulator (LQR), an important problem in control theory which can be thought of as an MDP optimization problem with an uncountably infinite state-space and average reward. The paper proves that in this setting, the naive policy gradient algorithm converges to the optimal policy, and the NPG algorithm converges with a faster rate guarantee. Kunnumkal and Topaloglu (Kunnumkal and Topaloglu, 2008) study the base-stock inventory control problem, and prove that an algorithm using the policy gradient framework achieves convergence to the optimal policy, with a bound on the convergence rate.
Follow-up study of these settings has recently demonstrated that these settings exhibit additional structure which allows these convergence results to hold (Bhandari and Russo, 2024). This additional structure relates to the interplay between policy gradient algorithms, which the above results focus on, and policy improvement algorithms. In particular, it has been shown that if the standard policy improvement algorithm exhibits no suboptimal stationary points, then policy gradient algorithms will track the standard policy iteration algorithm and thereby converge to the global optimum. This mechanism is the root cause behind the prior convergence results in the LQR and base-stock settings.
In contrast, such properties do not hold for general infinite-state MDPs, such as the queueing MDPs we study, which have much less structure than these previously-studied settings.
@@ 2.4. Queueing and Reinforcement Learning
In this paper, we study MDP optimization for a general class of queueing MDPs. A variety of papers have studied applying reinforcement learning to queueing problems, including using learning algorithms which build off of the NPG algorithm.
Dai and Gluzman (Dai and Gluzman, 2022) study variance reduction techniques for the PPO algorithm in the context of queueing models, and demonstrate empirically strong performance, converging towards an optimal policy in several queueing settings. The theoretical results in the paper focuses on reducing noise that arises from the sampling process, while our results focus on achieving convergence even in a setting where no noise is present at all.
Che et al. (Che et al., 2023) further employ a differentiable simulation-based modeling technique, allowing additional approximate gradient information to be derived, which is then used to improve the empirical reinforcement learning performance in a queueing context.
Wei et al. (Wei et al., 2023) have demonstrated that a “sample augmentation” technique can reduce the amount of sample data necessary to converge towards an optimal policy for pseudo-stochastic reinforcement learning settings, including queueing models. However, the paper focuses on discounted-cost finite-state queueing MDPs, rather than the infinite-state average-reward queueing models which are preferred elsewhere in queueing theory.
Adler and Subramanian (Adler and Subramanian, 2023) study an infinite-state, average-reward queueing MDP optimization setting, where the underlying queueing MDP is parameterized by an unknown parameter $\theta$. Queueing models with unknown dynamics are used as the motivating models. The paper assumes that an optimal policy $\pi_{\theta}^{*}$ is known for the MDP under any given parameter $\theta$, and focuses on learning the true parameter value $\theta^{*}$, and on minimizing the regret experienced while learning that parameter. Our result complements this paper, as we show that the NPG algorithm can be used to find optimal policies for MDPs in a setting where the underlying parameters are exactly known, providing the necessary input $\pi_{\theta}^{*}$ for the above result.
In work currently under submission, Chen et al. (Chen et al., ) study a primal-dual optimization technique for optimizing constrained MDPs, such as those encountered in queueing problems. They consider a Langrangian relaxation of the original MDP, and simultaneously optimize the policy and the dual weights of the relaxation. In the special case of an unconstrained MDP, the underlying primal algorithm they study is the NPG algorithm.
@@
2.5. Concurrent Work
In concurrent work currently under submission, Murthy et al. (Murthy et al., 2024) study a distinct, but related setting using the same framework as in this paper. In this paper, we study queueing MDPs in a noise-free setting, where the NPG updates can be computed exactly. Murthy et al. (Murthy et al., 2024) maintain the same queueing MDPs and generalize the setting to allow noisy NPG updates, but examine a restricted policy space, where all policies in the space are “uniformly stable”: All policies satisfy the Foster-Lyapunov theorem (Foster, 1953) with the same test function $V(s)$, the same bounded set, and the same drift $\epsilon$. Neither result is a strengthening of the other: This paper examines a general policy space under exact NPG updates, while the Murthy et al. (Murthy et al., 2024) paper allows noisy NPG updates but requires a restricted policy space with uniform stability.
@@ 3. Model, Assumptions, and Algorithm
In this paper, we study the Natural Policy Gradient (NPG) algorithm for MDP optimization, focusing on the MDPs corresponding to a highly general class of queueing models called the Generalized Switch with Static Environment (GSSE). We define general Markov Decision Process notation in Section 3.1. We define the NPG algorithm in Section 3.2. We define the GSSE class of queueing models in Section 3.3.
In Section 4, we show that the Natural Policy Gradient algorithm converges to an optimal policy for any queueing setting in the GSSE class, when initialized with the well-studied MaxWeight policy, which we define in Section 3.4. Our analysis of the NPG algorithm generalizes to queueing MDPs beyond the GSSE class, as well as beyond the queueing setting, to prove convergence for any infinite-state average-reward MDP that satisfies certain structural assumptions, when initialized with a sufficiently good initial policy. We specify these structural assumptions in Section 3.5.
@@ 3.1. Markov Decision Processes
The reward is given by a reward function $r(s,a)$ which maps a state-action pair to a real-valued reward. We likewise write $r(s,\pi)$ to denote the expected single-step reward associated with policy $\pi$ and state $s$. We define $r_{\max}(s)$ to be the maximum reward achievable in state $s$, $r_{\max}(s):=\max_{a}r(s,a)$. Note that we focus on reward-maximization rather than cost-minimization. Our primary reward function is the negation of the total queue length.
@@
(1) $J_{\pi}+V_{\pi}(s)=r(s,\pi)+\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{\pi}(s)}[V_{\pi}(s^{\prime})]$
@@
(2) $J_{\pi}+Q_{\pi}(s,a)$ $=r(s,a)+\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{a}(s)}[Q_{\pi}(s^{\prime},\pi)].$
To uniquely specify the additive constant $C$, we will adopt the convention that $V_{\pi}(\vec{0})=0$, for a specially identified state $\vec{0}$ defined in Assumption 1. In the queueing setting discussed in Section 3.3, $\vec{0}$ is the state with no jobs in the system.
Often, we are interested in the case where the action $a$ is sampled from a policy $\pi$’s action distribution $\pi(a\mid s)$. We define corresponding notation:
$Q_{\pi}(s,\pi^{\prime}):=\mathbb{E}_{a\sim\pi^{\prime}(\cdot\mid s)}[Q_{\pi}(s,a)]$
@@ 3.2. NPG Algorithm
3.3. Definition of Generalized Switch with Static Environment
The Generalized Switch with Static Environment (GSSE) is a queueing model with $n$ classes of jobs, and the state of the system is a size-$n$ vector $q_{\pi}(t)=\{q^{(i)}_{\pi}(t)\}$ specifying the number of jobs of each class present in the system under policy $\pi$. This model is a special case of the Generalized Switch model (Stolyar, 2004). Let $q_{\pi}=\{q^{(i)}\}$ denote the corresponding stationary random variable. Note that we use lower case $q$ to denote queue lengths and upper case $Q$ to denote state-action relative value.
There are $m$ service options. At each time step, prior to new jobs arriving, the scheduling policy selects a service option $j$. If the policy selects service option $j$, then $w_{i}^{j}$ jobs of class $i$ complete, where $w_{i}^{j}$ is a bounded, i.i.d. random variable. Jobs of each class complete independently. If on some time step $t$, $w_{i}^{j}(t)$ exceeds $q^{(i)}(t)$, the number of jobs of class $i$ present in the system, then all of the class $i$ jobs complete.
Our primary reward function $r(s,\cdot)$ of state $s=\{q_{i}\}$ under any action is the negative total queue length $r(s,\cdot)=-\sum_{i}q_{i}=-|q|_{1}$. We will also consider the setting where the reward is the negative $\alpha$-th moment of the total queue length, $-(\sum_{i}q_{i})^{\alpha}$, for a generic $\alpha\geq 1$.
Note that the assumption of a static environment can be relaxed. The static environment assumption is only used to prove drift properties of the MaxWeight policy in Lemma 11, which can be proven, with more effort, from more relaxed conditions on the environment.
3.4. MaxWeight
An important scheduling policy in the Generalized Switch model is the MaxWeight policy (Tassiulas and Ephremides, 1990), which chooses the service option which maximizes the inner product of the queue lengths and the service rates:
(3) $\text{MaxWeight}(q):=\arg\max_{j}\sum\nolimits_{i}q_{i}\mu^{j}_{i}$
In particular, MaxWeight is known to be throughput optimal, meaning that it keeps the system stable for all arrival rates within the capacity region (Stolyar, 2004; Hurtado Lange and Maguluri, 2022). Moreover, MaxWeight asymptotically minimizes the total queue length $\sum_{i}q_{i}$ in the heavy traffic limit (Stolyar, 2004) in our discrete-time setting.
In this paper, we use the MaxWeight policy to initialize our NPG algorithm as $\pi_{0}$.
3.5. Assumptions
The goal of this paper is to study queueing MDPs. We therefore define a general class of MDPs which include a wide variety of queueing MDPs. The class of MDPs studied in this paper are MDPs which satisfy two assumptions on the structure of our MDP: One on the rewards in the MDP, and one on the connectedness of high-reward states. We verify these assumptions for GSSE queueing MDPs in Section 5.8.
First, we make mild assumptions on the reward structure of the MDP, which are straightforwardly satisfied by a wide variety of queueing MDPs:
Assumption 1 (Reward structure).
  1. (a)
  1. (b)
  1. (c)
  2. (d)
In the queueing setting, reward is $r(s,a)=-|q|_{1}$, $c_{\max}=0$, and $\vec{0}$ is the empty-queue state.
Second, we assume that the high-reward states are uniformly connected, under an arbitrary policy. For a queueing MDP, this states that it’s possible to move from any short-queue state to any other short-queue state in a uniformly bounded number of steps under an arbitrary policy.
Assumption 2 (Uniform connectedness of high-reward states).